An amplifier has an input signal of 5V at 0.1mA. The output of the amplifier provides 10mA to a 450Ω load
determine : a) the current , voltage and power gains
b) the input resistance to the amplifier
1
Expert's answer
2020-10-16T10:59:51-0400

Explanations & Calculations


  • Input current = 0.1mA
  • Output current = 10mA
  • Output voltage = "\small 10mA \times450 \Omega = 4.5V"


  • Therefore, current gain = "\large \frac{10mA}{0.1mA} \small = \bold{100}"
  • Voltage gain = "\large \frac{4.5V}{5V} \small = \bold{0.9}"


  • Power gain = "\large \frac{\text{output power}}{\text{input power}} \small = \large \frac{4.5V\times 10mA }{5V \times 0.1mA} \small =\bold{90}"


  • Input resistance = "\large \frac{\text{input voltage}}{\text{input current}} = \frac{5V}{0.1mA} \small =\bold{50k \Omega}"